OpenStudy (anonymous):

How do you get from x^4-6x^2+8 to (x^2-2)(x-2)(x+2)? Is there a formula?

OpenStudy (mathstudent55):

You need to factor. When you look at x^4 - 6x^2 + 8, you may notice that if you substitute u for x^2, by letting u = x^2, then u^2 = x^4. Then you have: u^2 - 6u + 8 Now this looks like a simple 2nd degree polynomial that should be easy to factor: (u - 4)(u - 2) Do you follow so far?

OpenStudy (mathstudent55):

Once you have (u - 4)(u - 2), substitute back x^2 = u (x^2 - 4)(x^2 - 2) The left factor is the diference of two squares, so we can factor it: (x+ 2)(x - 2)(x^2 - 2)

OpenStudy (anonymous):

OMG, I can't believe I didn't get this on my own. Thanks for your help mathstudent55, I should of figured it is something simple like this.